# AR-KAN: Autoregressive-Weight-Enhanced Kolmogorov–Arnold Network for Time Series Forecasting

### v1

@@ Abstract
[…]
@@ I Introduction
Our experiments confirm the effectiveness and generalizability of AR-KAN. On almost-periodic functions, it matches the performance of ARIMA. Moreover, across 18 real-world datasets from diverse domains, AR-KAN outperforms all baselines and achieves the best results on 72% of them (13 out of 18). These results highlight AR-KAN’s robustness, adaptability, and promise as a unified framework for time series forecasting.
@@
Section II introduces the background, including time series forecasting tasks, ARIMA, MLP, and KAN models. Section III presents the Universal Myopic Mapping Theorem and explains how it inspires the overall architecture of our AR-KAN model. Section IV describes the experiments conducted to demonstrate the effectiveness and generalizability of AR-KAN, including evaluations on two constructed almost-periodic functions and 18 real-world time series. Finally, Section V concludes the paper and discusses potential directions for future research.
@@ IV Experiments
We conduct experiments in two parts to demonstrate both the effectiveness and generalizability of AR-KAN. First, we perform experiments on noisy almost-periodic functions to show that modern models fall short of traditional ARIMA models in terms of spectral analysis, while our AR-KAN achieves performance comparable to ARIMA. Then, we extend the evaluation to 18 real-world datasets from Rdatasets[40], demonstrating that AR-KAN achieves the best performance on 72% of them. The detailed experimental settings are provided in appendices.
TABLE II: Test loss (MSE) of various models on Rdatasets
Datasets ARIMA AR-KAN AR-MLP KAN MLP LSTM FAN FNO LLMTime
a10_ts 0.1441 0.1353 0.4775 2.5033 2.2638 0.8809 0.4913 0.3851 0.3457
airpass_ts 0.3329 0.0706 0.0871 0.3046 0.3025 0.4249 0.5163 0.6982 0.1937
ausbeer_ts 0.0418 0.0357 0.0741 0.1031 0.5102 0.0692 0.0802 0.1114 0.0436
auscafe_ts 0.3301 0.3813 0.1180 2.6312 0.7564 0.3746 1.4820 3.4769 0.4463
BJsales_ts 0.3241 0.0032 0.0261 0.0358 0.7849 0.0643 0.2393 1.0370 0.0131
bricksq_ts 0.2080 0.0502 0.0823 0.2542 0.2769 0.2961 0.9625 0.2607 0.2541
co2_ts 0.0218 0.0014 0.0064 0.3079 0.0460 0.1640 0.1584 0.1963 0.0109
discoveries_ts 1.6030 2.1695 2.3091 1.7269 1.6264 1.7949 1.0153 0.8469 1.2922
economics_df_ts 3.5659 0.0845 0.4398 2.3047 1.8717 2.2520 1.8607 7.7670 1.6490
elec_ts 0.2731 0.0069 0.0060 0.093 0.0436 0.1833 0.3258 0.0625 0.0727
elecdaily_mts 0.4331 0.2123 0.2573 0.441 0.5366 0.6792 0.8981 0.5919 0.6127
elecequip_ts 0.3159 0.1528 0.1346 0.5968 0.5538 0.8984 0.7761 0.4870 1.4010
euretail_ts 0.4967 0.9964 1.3328 1.1984 0.4226 1.6740 1.1469 0.1821 1.5009
goog200_ts 4.7135 0.1228 0.8096 3.6632 3.5888 3.2584 3.0580 7.9012 1.1351
gtemp_both_ts 2.2374 0.2936 0.5328 3.2225 2.2946 1.6629 2.8678 1.6660 5.0938
h02_ts 0.2726 0.1263 0.1782 1.3708 0.5258 0.2103 0.6003 0.8209 0.1371
hsales2_ts 0.5781 0.5232 0.6301 2.1212 0.8286 1.7787 1.7065 0.8850 0.5667
hyndsight_ts 0.8729 0.2471 0.3961 1.5734 0.4892 0.5929 0.6164 0.6793 0.4510
Note: Bold numbers indicate the minimum value in each row.
Figure 4: Performance of ARIMA, AR-KAN and FNO on cases of Rdatasets (left: a10_ts, right: ausbeer_ts).
@@ IV-B Rdatasets
To further demonstrate the generalizability of AR-KAN, we select 18 time series from the Rdatasets for evaluation, including a10_ts, airpass_ts, and others. These time series span a wide range of domains and exhibit diverse temporal characteristics. The performance results across all models are summarized in Table II. Typically, the outcomes of certain series (a10_ts and ausbeer_ts) produced by ARIMA, AR-KAN, and FNO are shown in Fig. 4.
The results clearly demonstrate the superior generalizability of AR-KAN across a diverse set of real-world time series. AR-KAN achieves the best performance on 13 out of 18 datasets (72%), demonstrating a notable superiority compared to the other baseline models. This consistent advantage across datasets of varying domain, scale, and statistical properties underscores AR-KAN’s capacity to adapt effectively to a wide range of temporal dynamics.
In contrast, ARIMA, despite its strong performance on the synthetic almost-periodic task, performs best on only a single real-world dataset. This sharp performance drop highlights ARIMA’s limited expressiveness when faced with the complexity and noise inherent in practical scenarios. Similarly, Fourier-based models such as FAN and FNO, though designed to leverage spectral structures, fail to generalize effectively. Their reliance on strong spectral priors appears insufficient for capturing the heterogeneous and often non-stationary nature of real-world signals.
Other deep learning models such as MLP, LSTM, and even KAN alone exhibit fluctuating performance and lack robustness across datasets. These models either overfit to local patterns or underperform due to inadequate inductive biases. By the way, with the rapid advancement of large language models (LLMs), the performance of LLMTime[42] has improved from being consistently inferior to ARIMA[41] to roughly on par with it. However, it still lags significantly behind AR-KAN. LLMs still have a long way to go in the domain of time series modeling.
By contrast, AR-KAN offers a principled blend of classical signal modeling and modern neural adaptation. It not only inherits the frequency-awareness of traditional methods, but also leverages the expressive power of neural architectures to model nonlinear and non-stationary behaviors. This versatility positions AR-KAN as a highly effective solution for a wide spectrum of time series modeling tasks.
@@ V Conclusion
In this paper, we identified the limitations of existing neural networks in spectral analysis and demonstrated through experiments that they underperform traditional ARIMA models when dealing with almost-periodic functions. Motivated by this observation, and based in the Universal Myopic Mapping Theorem, we proposed AR-KAN, which is a novel model that effectively integrates the autoregressive component of ARIMA with the expressive nonlinear modeling capability of KAN. Experimental results show that AR-KAN achieves performance comparable to ARIMA on almost-periodic functions, and outperforms all baselines on 13 out of 18 real-world datasets, demonstrating both its effectiveness and strong generalizability.
More broadly, the contribution of this paper goes beyond AR-KAN as a single model; it brings renewed attention to Universal Myopic Mapping Theorem. This framework enables the integration of classical signal processing techniques with modern neural architectures, thereby offering interpretability and adaptability, and improving the versatility of neural models in time series tasks. We hope that future research will further explore this direction and extend the framework to more complex and diverse time series analysis scenarios.

### v2

# AR-KAN: Autoregressive-Weight-Enhanced Kolmogorov-Arnold Network for Time Series Forecasting

@@ Abstract
Traditional neural networks struggle to capture the spectral structure of complex signals. Fourier neural networks (FNNs) attempt to address this by embedding Fourier series components, yet many real-world signals are almost-periodic with non-commensurate frequencies, posing additional challenges. Building on prior work[41] showing that ARIMA outperforms large language models (LLMs) for forecasting, we extend the comparison to neural predictors and find ARIMA still superior. We therefore propose the Autoregressive-Weight-Enhanced Kolmogorov–Arnold Network (AR-KAN), which integrates a pre-trained AR module for temporal memory with a KAN for nonlinear representation. The AR module preserves essential temporal features while reducing redundancy. Experiments demonstrate that AR-KAN matches ARIMA on almost-periodic functions and achieves the best results on 72% of Rdatasets series, with a clear advantage on data with periodic structure. These results highlight AR-KAN as a robust and effective framework for time series forecasting. Our code is available at https://github.com/ChenZeng001/AR-KAN.
@@ I Introduction
Extensive experiments demonstrate AR-KAN’s effectiveness. On almost-periodic functions, it matches ARIMA’s performance. And on real-world datasets from Rdatasets[40], it outperforms baselines on 72% of tasks. The few cases where AR-KAN does not dominate correspond mostly to datasets with weak or nearly absent periodicity, which are inherently difficult to forecast without additional prior information. When focusing on datasets with stronger periodic components, AR-KAN shows a clear and consistent advantage, highlighting its robustness, adaptability, and potential as a unified framework for time series forecasting.
@@
Section II introduces the background, including time series forecasting tasks, ARIMA, MLP, and KAN models. Section III presents the Universal Myopic Mapping Theorem and explains how it inspires the overall architecture of our AR-KAN model. Section IV describes the experiments conducted to demonstrate the effectiveness and generalizability of AR-KAN, including evaluations on two constructed almost-periodic functions and real-world time series. Finally, Section V concludes the paper.
@@ IV Experiments
We conduct experiments in two parts to demonstrate both the effectiveness and generalizability of AR-KAN. First, we perform experiments on noisy almost-periodic functions to show that modern models fall short of traditional ARIMA models in terms of spectral analysis, while our AR-KAN achieves performance comparable to ARIMA. Then, we extend the evaluation to real-world datasets from Rdatasets[40], and observe that AR-KAN achieves the best performance on 72% of datasets. The few cases where it underperforms mostly involve weak or absent periodicity, which are inherently hard to predict without prior knowledge. For strongly periodic data, AR-KAN shows a clear and consistent advantage.
TABLE II: Test loss (MSE) of various models on Rdatasets with periodicity strength
Datasets Periodicity Strength ARIMA AR-KAN AR-MLP KAN MLP LSTM FAN FNO LLMTime
airpass_ts 41.28% 0.3329 0.0706 0.0871 0.3046 0.3025 0.4249 0.5163 0.6982 0.1937
goog200_ts 34.24% 4.7135 0.1228 0.8096 3.6632 3.5888 3.2584 3.0580 7.9012 1.1351
euretail_ts 18.39% 0.4967 0.9964 1.3328 1.1984 0.4226 1.6740 1.1469 0.1821 1.5009
ausbeer_ts 14.63% 0.0418 0.0357 0.0741 0.1031 0.5102 0.0692 0.0802 0.1114 0.0436
h02_ts 13.31% 0.2726 0.1263 0.1782 1.3708 0.5258 0.2103 0.6003 0.8209 0.1371
BJsales_ts 10.08% 0.3241 0.0032 0.0261 0.0358 0.7849 0.0643 0.2393 1.0370 0.0131
a10_ts 6.36% 0.1441 0.1353 0.4775 2.5033 2.2638 0.8809 0.4913 0.3851 0.3457
hsales2_ts 4.20% 0.5781 0.5232 0.6301 2.1212 0.8286 1.7787 1.7065 0.8850 0.5667
co2_ts 3.69% 0.0218 0.0014 0.0064 0.3079 0.0460 0.1640 0.1584 0.1963 0.0109
hyndsight_ts 2.61% 0.8729 0.2471 0.3961 1.5734 0.4892 0.5929 0.6164 0.6793 0.4510
bricksq_ts 1.33% 0.2080 0.0502 0.0823 0.2542 0.2769 0.2961 0.9625 0.2607 0.2541
elecequip_ts 1.25% 0.3159 0.1528 0.1346 0.5968 0.5538 0.8984 0.7761 0.4870 1.4010
elecdaily_mts 1.16% 0.4331 0.2123 0.2573 0.4410 0.5366 0.6792 0.8981 0.5919 0.6127
gtemp_both_ts 0.50% 2.2374 0.2936 0.5328 3.2225 2.2946 1.6629 2.8678 1.6660 5.0938
discoveries_ts 0.22% 1.6030 2.1695 2.3091 1.7269 1.6264 1.7949 1.0153 0.8469 1.2922
elec_ts 0.16% 0.2731 0.0069 0.0060 0.0930 0.0436 0.1833 0.3258 0.0625 0.0727
economics_df_ts 0.00% 3.5659 0.0845 0.4398 2.3047 1.8717 2.2520 1.8607 7.7670 1.6490
auscafe_ts 0.00% 0.3301 0.3813 0.1180 2.6312 0.7564 0.3746 1.4820 3.4769 0.4463
Note: Bold numbers indicate the minimum value in each row. Rows are shaded in red when periodicity is present ($\geq 0.5\%$) and in blue when periodicity is weak ($0.5\%$).
Figure 4: Performance of ARIMA and AR-KAN on two different types of time series in Rdatasets. the left column shows results on a10_ts (a highly periodic series), and the right column shows results on discoveries_ts (a nearly non-periodic series).
@@ IV-B Rdatasets
We further evaluate AR-KAN on real-world series from Rdatasets[40]. For each dataset, we quantify its Periodicity Strength as the ratio of the energy of the seasonal component (obtained by STL decomposition[43]) to the total energy of the original series:
$\text{Periodicity Strength}=\frac{\|x_{\text{seasonal}}\|_{2}^{2}}{\|x\|_{2}^{2}}.$
The period used in STL decomposition is determined by the largest nonzero-lag peak in the series’ autocorrelation function. Our results in TABLE II show that AR-KAN achieves the best performance on 72% of datasets. Moreover, its advantage grows with increasing Periodicity Strength, indicating that AR-KAN excels on series with clear or strong periodic components. In contrast, its few underperforming cases correspond to datasets with weak or nearly absent periodicity, which are inherently hard to forecast without additional prior knowledge and remain challenging for all competing methods. They are illustrated in Fig. 4. By the way, Fourier-based models fail to generalize despite their spectral priors, and other neural baselines show similar limitations.
LLM-based methods such as LLMTime[42] approach ARIMA[41] in performance but still fall short of AR-KAN, indicating that LLMs are not yet mature for time series forecasting. Both LLMs and neural networks struggle to capture frequency structure, whereas AR offers a statistical view for genuine spectral analysis. By combining this frequency-aware modeling with KAN’s nonlinear expressivity, AR-KAN achieves robust, domain-general performance.
@@ V Conclusion
In this paper, we reveal that existing neural networks struggle with spectral analysis and often underperform ARIMA on almost-periodic functions. Guided by the Universal Myopic Mapping Theorem, we propose AR-KAN, which combines ARIMA’s autoregressive memory with KAN’s nonlinear expressivity. Experiments show that AR-KAN matches ARIMA on almost-periodic functions and outperforms baselines on 72% of real-world datasets, with its advantage growing on series with clear periodic patterns. These results highlight AR-KAN as a robust and unified framework for time series forecasting.